OpenStudy (anonymous):

Algebra 2 help? Please? >.<

OpenStudy (anonymous):

how can i help if there is no question?

OpenStudy (anonymous):

I was seeing who would respond. I'm kinda new to this.

OpenStudy (anonymous):

that's okay, but it's best practice to just ask the question and see if anyone answers. If not you can always bump it.

OpenStudy (anonymous):

Launch Area:___(1, 2)___ Point A:___(0, 3)___ Point B:___(-3, 0)___ Point C:___(-1, -4)___ Graph the coordinates of the specifics points in space your spacecraft will travel to. You must show your work on each question below. Determine the equation of the line, in standard form, that will get your spacecraft from the Launch Area to Point A, Ax+By=C Determine the equation of the line, in point-slope form, that will get your spacecraft from Point A to Point B. y - y1 = m(x - x1) Determine the equation of the line, in slope-intercept form, that will get your spacecraft from Point B to Point C. y=mx+b Convert the equation you arrived at in question 2 into slope-intercept form. Make sure to include all of your work. y=mx+b Reflect back on this scenario and each equation you created. Would any restrictions apply to the domain and range of those equations? Explain your reasoning using complete sentences. So far, I think I have the correct forms, I just don't know how to put them into affect.

OpenStudy (anonymous):

or tag people you know

OpenStudy (anonymous):

that seems rather logical xD

OpenStudy (anonymous):

Here's the graph that's supposed to be included

OpenStudy (anonymous):

so, the first part of the question is finding an equation to move from (1,2) to (0,3)

OpenStudy (anonymous):

i'm not sure what type of file that is supposed to be, i can't open it and i don't recognise it as any known filetype.

OpenStudy (anonymous):

what the question says is to use standard form and I believe that's what it is, but how would I use that to get from (1,2) to (0,3)?

OpenStudy (anonymous):

I apologize, it's saved in an online format, let me get it to a png.

OpenStudy (anonymous):

nvm i managed to open it

OpenStudy (anonymous):

hmm let me think, i'm not that used to standard form

OpenStudy (anonymous):

Take your time man, I'll have all week to have this done xD

OpenStudy (anonymous):

alright, give me a few minutes to read up on the subject

OpenStudy (anonymous):

alright, apparently one way to do it is to get the slope intercept form and than manipulate it into the standard form.

OpenStudy (anonymous):

First, let's get the slope. Which you can get by answering the question: "when I move 1 step in the x direction, what will happen to the y direction?"

OpenStudy (anonymous):

since we start from (1,2) and move to (0,3). So we move -1 space in the x direction and +1 space in the y direction we can conclude our slope must be -1. (for every step in the x direction we go -1 step in the y direction)

OpenStudy (anonymous):

so the slope intercept form is y=-x+b and the b can be found by inserting actual coordinates into the equation. (1,2) is the coordinate of the launch pad, so we can insert that. 2 = -1 + b so b = 3 so the slope intercept form is y = -x + 3

OpenStudy (anonymous):

rewriting it to standard form it's y+x=3 in this case it's as simple as moving the x to the other side and flipping the sign. Sometimes you have to multiply the whole equation to get rid of fractions

OpenStudy (anonymous):

@alanate did i went to fast or do you get it?

OpenStudy (anonymous):

I think I get it a little more than what I have from the lesson, thank you c: